a soap bubble of radius 3 cm is charged with 9 nC . the excess pressure inside the bubble , if surface tension of soap solution is 3× 10-3 N/m.ans is .36 N/m2

Grade:12

2 Answers

Arun
3 years ago
dear stdent
excess pressre due to surface tension =4S/r
here S=3*10−3 r= 3cm excess pressre = 4SR=4*3*10−3/3*10−2 = 0.4now excess ressure due to electric field = σ22ε=q22A2εq22(4πr2)ε here ε=8.85*10−12 here q= 9nm=9*10−9C r=3cm =3*10−2putting all values we get pressure due to electric field= 0.035N/m2its is outward to the spherical bubble so net pressure= 0.4− 0.035 =0.36N/m2
Vikas TU
3 years ago
dear stdent
excess pressre due  to  surface tension =4S/r
here  S=3*10^−3 r= 3cm excess pressre = 4S/R
=4*3*10^−3/3*10^−2 
= 0.4now excess ressure due to electric field 
= σ^2/2ε=q^2/2A^2ε
q^2/2(4πr^2)ε 
here ε=8.85*10^−12 
here q= 9nm=9*10^−9C 
r=3cm =3*10−2putting all values we get pressure due to electric field= 0.035N/m2
its is outward to the spherical bubble so net pressure= 0.4− 0.035 =0.36N/m2